How do I explain finding this answer to my 4th grader? At a department store a hat and a shirt cost $22. The hat costs $12 less than the shirt. How much does the hat cost at the store? How much does the shirt cost?
Answer by Alan3354(69443)

At a department store a hat and a shirt cost $22. The hat costs $12 less than the shirt. How much does the hat cost at the store? How much does the shirt cost?
----------------
h + s = 22
h = s - 12
----------
sub for h in the 1st equ
(s - 12) + s = 22
2s - 12 = 22
2s = 34
s = 17
h = 5